Use the z-table to find the area under the normal curve that is located between z=0 and each of the following.

AREA

1.z=0.56

2.z=1.32

3.z=1.05

4.z=-2.18

5.z=2.58


1
Expert's answer
2021-03-15T08:30:17-0400

S = f(z) - 0.5

1. f(0.56) = 0.71226, S = 0.71226 - 0.5 = 0.21226

2. f(1.32) = 0.90658, S = 0.90658 - 0.5 = 0.40658

3. f(1.05) = 0.85314, S = 0.85314 - 0.5 = 0.35314 

4. f(-2.18) = 0.01463, S = 0.5 - 0.01463 = 0.48537

5. f(2.58) = 0.99506, S = 0.99506  - 0.5 = 0.49506